# On the Power Saving in High-Speed Ethernet-based Networks for Supercomputers and Data Centers

### v2

@@ 4 Evaluation
In this section, we will evaluate the mechanisms previously proposed, which we have implemented in SAURON [sauron], our OmNet-based network simulator. Using this simulator, we have prepared some experiments to test power-saving strategies. Regarding the network traffic, our simulator employs the VEF-TraceLib [VEF-traces, Andujar16JSC, andujar_extending_2023] library to inject traces from real applications into the simulator. TraceLib also provides us with a statistic of CPU usage, which will help us model the dynamic CPU usage-to-power conversion. We assume that the power consumption for nodes is mostly determined by the CPU, whose power is a direct function of usage, with a minimum power draw at idle. We also use the TopGen [topgen] library and its file generator to model large-scale networks with multiple topologies.
@@
The scenario we have chosen for the experiments is a Megafly [flajslik_megafly_2018] topology composed of 4160 computing nodes. The network is arranged in 65 groups of 64 nodes, each group employing 16 switches with a radix of 16 (1040 switches in total). The inter-group connectivity is as described in the original Megafly publication, with one global link containing every pair of groups. Our decision on this topology is its scalability, flexibility, and popularity in HPC environments [mflyadapt]. This topology can exploit the relative locality of nodes located in the same group for fast efficient communication in disjoint network partitions that exchange their messages through high-bandwidth global channels. This node arrangement can help perform partial calculations in the same-group nodes and use the global links for reductions. AI workloads, and most recently, Large Language Models (LLMs), can also benefit from this topology, as several kinds of parallelization employ different broadcast domains. Compared to the Dragonfly, the main advantages are the use of low-radix switches and the inherent absence of deadlocks. The routing algorithm employed is deterministic and minimal. When a packet has its source and destination in the same group, the algorithm behaves like D-mod-k [PGFT]. However, when the destination is in another group, the packet travels to the spine switch that will lead to the destination group through a global link. When the packet arrives at the destination group, the downward path is the same as in D-mod-k, with only one possible route.
Table 5 shows the power consumption of the different elements that make up the whole system. These figures are based on information on modern server processors and expected network power draw provided by Eviden. The asterisk on minimum link power consumption depends on the different power-saving states that are employed. The total link number accounts for the link on every switch times the radix, with the addition of every NIC. The system is considered idle when the nodes are not doing any computation and thus, consuming their minimum power. Switch power only comprises the device power draw with no links connected.
@@
Min Max Min Max
Switch 250 W 250 W 1040 260 KW 260 KW
Node 800 W 1200 W 4160 3.328 MW 4.992 MW
Link * 24 W 20800 * 499.2 KW
* 5.751 MW
Table 5: Power consumption of the different components of the system.
On the contrary, under full load, the system consumes approximately 3.788 MW. In both cases, the contribution to power consumption by the network is the power consumption of switches and links, 260 KW and 124.8 KW, respectively, adding up to 759.2 KW. In proportion, the network accounts for 12.214 % of power when the system is idling, and up to 8.68 % under full load when all nodes are computing. A way to save power on links is to employ the EEE standard and its power-saving link states.
In Table 6, we define the power parameters for each power-saving state, such as the power reduction and the time to transition from the power-saving state to the active state ($t_{w}$), and its counterpart ($t_{s}$), derived from EEE (which we showed in Table 3). The table also shows the percentage of power that the system devotes to the network upon applying those values to our selected scenario. The calculi for the power contribution of different elements is similar to the one by [ANDUJAR2023126].
Port power state Power $t_{w}$ $t_{s}$ Network/total power Links/network power
@@
Wake 24 W - - 18.575 % 13.201 % 12.214 % 8.68 %
Fast Wake 9.6 W 375 ns 200 ns 12.136 % 8.432 % 5.272 % 8.68 %
Deep Sleep 2.4 W 4.48 µs 2 µs 8.519 % 5.845 % 1.372 % 8.68 %
@@
For the evaluation, we have selected traces of two HPC applications (LAMMPS, PATMOS) and two distributed Deep Learning training sessions (MLWF [MLWF] and AlexNet [ALEXNET]) to study the network behavior in both contexts. For each application, we will first describe the behavior of the application and the potential for energy saving. First, we will compare how much energy we save with different techniques in comparison to not using any power-saving technique. Non-energy metrics include the degradation that is introduced by power-saving strategies. Mainly, we want to focus on the increase in execution time and latency. Specifically, the latency metric we are exploring is the packet latency, which measures the time from the packet generation to its arrival at the destination.
@@ 4.1 Evaluation for the LAMMPS application
In summary, LAMMPS creates a mix of localized and global communication patterns, with MPI_AllReduce identified as the primary contributor to potential network bottlenecks, especially in large-scale simulations. Being an all-to-all communication operation, MPI_AllReduce will use most paths between nodes. This means that this operation will have the most degradation if all links are in a power-saving state. Figure 6 shows the network efficiency or throughput (normalized against the total network bandwidth) for the complete execution time (2.431 seconds) in the configured system. The total energy consumed is 9.993 MJ by nodes and 1.846 MJ by the network, totaling 11.839 MJ for the whole system. We can see that after a second of computation, the network activity starts being intermittent with several traffic spikes due to the inter-node collective operations.
@@ 4.1.1 Results with fixed $t_{PDT}$ values
(b) Total energy savings.
@@
In general, when using a fixed $t_{PDT}$ value, we can see that the maximum energy savings are close to 10% when a LAMMPS application is run using our power model and the $t_{PDT}$ values are higher than or equal to 100 µs. With this $t_{PDT}$, the overheads in execution time and latency are also negligible. Fast Wake is more sensitive to smaller $t_{PDT}$ values and achieves better energy savings with smaller $t_{PDT}$ values. Finally, note that with fixed $t_{PDT}$ values larger than 1msms there are barely energy savings.
@@ 4.1.2 Results using PerfBound and PerfBoundCorrect
The overhead on execution time has a direct impact on the energy consumed, as seen in Figure 8(b). Both PerfBound and PerfBoundCorrect provide similar levels of energy savings, with slightly higher values using PerfBoundCorrect. While PerfBound with Deep Sleep does not lead to energy savings, for thresholds of 1 % and 2 % PerfBoundCorrect manages to revert that situation. When using a high threshold, however, energy saving does not occur because the calculated $t_{PDT}$ values are too short.
Finally, Figure 8(c) shows that PerfBoundCorrect, in all cases and with either power-saving state, reduces latency significantly. It is clear that, particularly when using Deep Sleep, a circular buffer to control the size of the histogram produces the highest overhead.
@@ 4.2 Evaluation for the PATMOS application
PATMOS [PATMOS] is a Monte Carlo neutron transport code developed at the CEA (France), designed to prototype applications in nuclear safety and radiation shielding.
@@
During each simulation cycle, particles are evenly distributed among threads, with each thread responsible for simulating particle histories and updating the associated scores. By default, a static scheduling is used so that the parallelization strategy is deterministic, ensuring reproducibility and simplifying debugging.
@@ 4.2.1 Results with fixed ${t}_{PDT}$ values
The first statistic we will analyze is the execution time, as it is the one that is most noticeable in terms of overhead, and a longer execution times have a negative impact on total energy consumption. In our experiments, applying any power-saving technique did not affect this application. The reason behind this is that performance overhead is cumulative with the number of packets that get stalled on output ports. Therefore, as the network is only used at the start and end of the simulation, the effect on execution time is virtually non-existent.
@@
(b) Total energy savings.
@@ 4.2.2 Results using PerfBound and PerfBoundCorrect
The execution time with any fixed $t_{PDT}$ value is nearly identical to the configuration to not applying any power-saving. PerfBound, focusing on minimizing the overhead on execution time, is no exception. Every execution time with any threshold or histogram management scheme yields the same results.
@@ 4.3 Evaluation for the MLWF application
The MLWF (Machine Learning Weather Forecast) [MLWF] application is developed within the context of the MAELSTROM European project. The trace was collected during a training session using Horovod [sergeev2018horovod2]. The traffic pattern in this application, in particular, and in Deep Learning training sessions in general, usually involves the usage of MPI collective operations for data exchange. Particularly, the most predominant traffic portion comes from AllReduce operations. Usually, every layer of the neural network that is being trained runs a repetitive sequence of Gather-Gatherv-Broadcast-Broadcast until the weights converge, then the AllReduce is invoked to have the information on every node [andujar_extending_2023].
The network efficiency results for this trace are shown in Figure 12. We have opted to show them on a logarithmic scale to illustrate that any traffic rate, regardless of the value, provides great difficulty in finding opportunities for power saving.
@@ 4.3.1 Results with fixed ${t}_{PDT}$ values
(c) Total energy savings.
@@
The results for energy saving for this application are shown in Figure 13(c). As shown, $t_{PDT}$ values of 10 and 100 µs increase the energy saved in both power-saving states. Lower values provide marginal power-saving values or even increased energy usage in that proportion. Indeed, even if the traffic between the large bursts at the start and end of the applications, there is communication that is not negligible for us. However, once we use higher values, the energy reduction begins to diminish. This is due to inactivity, 99 % of periods being within the millisecond range. This is also the reason why using a 1-second $t_{PDT}$ provides more power saving than a millisecond-scale one.
@@ 4.4.1 Results with fixed ${t}_{PDT}$ values
(c) Total energy savings.
@@ 5 Outcomes and future work
In this work, we have provided a detailed examination of the implications of power in interconnection networks. First, we explored the simplest technique by using fixed $t_{PDT}$ values, which gave us an upper bound for degradation. Later on, we have discussed more sophisticated approaches and solutions that have been proposed, focusing on HPC environments. After careful examination of our own implementation of the PerfBound technique and the results we gathered, we have found several key aspects that could be improved. As we have mentioned in earlier sections, the results are significant compared to the original PerfBound technique, and proves that giving more breathing room it can actually be beneficial to energy efficiency. Furthermore, we also show that other ostensibly trivial aspects like histogram management can have an impact on the energy saved, which was not examined in the original work.
@@
In our humble opinion, our solution is subject to improvements and several new questions arise. For example, whether providing our dynamic $t_{PDT}$ value calculation with some recency bias would improve the performance on very repetitive patterns.

### v3

@@ 4 Evaluation
In this section, we evaluate the power management mechanisms previously proposed, which we have modeled in our OMNeT++-based SAURON simulator [40]. First, we describe the experiment configurations and the network traffic workloads. Next, we include and analyze the experiment results for these workloads.
4.1 Experiments configuration
The SAURON simulator follows an event-driven approach and models the network devices (i.e., network interfaces and switches) at the packet level. We have assumed the network parameters shown in Table 6.
Parameter Value
Arbitration Aging (crossbar), Round-Robin (output port)
Link speed 400 Gbps
MTU 9600 Bytes
Network topology Megafly (see Section 4.2)
Node count 4160
PFC GO limit 3200 Bytes (6.51 % occupancy)
PFC STOP limit 37900 Bytes (77.11 % occupancy)
Queue size 49152 Bytes
Switch architecture CIOQ
Switch count 1040
Switch radix 16
Switch speed-up 2.44
Table 6: System configuration parameters for the experiments.
Regarding the network traffic, our simulator employs the open-source VEF Traces framework [4, 5, 3], which allows us to record MPI-based traffic in the VEF traces and use the TraceLib library to inject those messages into the SAURON simulator. The VEF traces have been collected in our own cluster CELLIA [27], mainly composed of multiprocessor nodes. As we describe later, we have selected traces from different applications, such as HPC and Deep Learning. The TraceLib library also provides information about the computing activity at system nodes, which helps us model their power consumption. Specifically, this power is a direct function of usage, with a minimum power draw at idle. Thus, the dynamic power consumption of a node ($P$) is mainly determined by its maximum and minimum power consumption, and the percentage of utilization based on computing activity (e.g., CPU or accelerator activity): $P=P_{min}+(P_{max}-P_{min})~\times\%~utilization$.
Regarding the network topology, the SAURON simulator uses the TopGen [39] library and its file generator to model large-scale networks with multiple topologies. Specifically, we have selected the Megafly topology for these experiments. As we discuss in the next subsection, Megafly is more power-efficient compared to other alternatives.
4.2 Discussion on the Topology selection
For the evaluation experiments, we assume a Megafly [16] topology (a.k.a., Dragonfly+ [34]) composed of 4160 computing nodes. The network is arranged in 65 groups of 64 nodes, each group employing 16 switches with a radix of 16 (1040 switches in total). The inter-group connectivity is as described in the original Megafly publication, with one global link containing every pair of groups. We have selected this topology for the experiments due to its scalability, flexibility, and growing popularity in HPC environments [25].
@@ 4.2 Discussion on the Topology selection
Specifically, Megafly topologies can exploit the relative locality of nodes placed in the same group for fast efficient communication in disjoint network partitions that exchange their messages through high-bandwidth global channels. This node arrangement can help perform partial calculations in the same-group nodes and use the global links for reductions. AI workloads, and most recently, Large Language Models (LLMs), can also benefit from this topology, as several kinds of parallelization employ different broadcast domains. Compared to the Dragonfly, the main advantages are the use of low-radix switches and the inherent absence of deadlocks. The routing algorithm modeled in the simulator is deterministic and minimal. When a packet has its source and destination in the same group, the algorithm behaves like D-mod-k [41]. However, when the destination is in another group, the packet travels to the spine switch that will lead to the destination group through a global link. When the packet arrives at the destination group, the downward path is the same as in D-mod-k, with only one possible route.
Regarding power consumption, Megafly is a power-efficient alternative compared to other options such as 3D-Tori or RLFTs. Table 7 shows a comparison in power consumption of the different devices of an HPC system (i.e., Switch, Node, and Links) for a Megafly, 3D-Torus, and RLFT networks using a similar number of node counts.
@@
Min Min Max
Megafly Switch 250 W 1040 260 KW (6.36 %) 260 KW (4.52 %)
Node 800 W - 1200 W 4160 3.328 MW (81.42 %) 4.992 MW (86.80 %)
Link 24 W 20800 499.2 KW (12.21 %) 499.2 KW (8.68 %)
4.087 MW 5.751 MW
Torus Switch 250 W 1000 250 KW (6.07 %) 250 KW (4.37 %)
Node 800 W - 1200 W 4000 3.2 MW (77.63 %) 4.800 MW (83.89 %)
Link 24 W 28000 672 KW (16.3 %) 672 KW (11.74 %)
4.122 MW 5.722 MW
RLFT Switch 250 W 845 211.250 KW (4.85 %) 211.250 KW (3.45 %)
Node 800 W - 1200 W 4394 3.515 MW (80.64 %) 5.273 MW (86.20 %)
Link 24 W 26364 632.736 KW (14.52 %) 632.736 KW (10.34 %)
4.359 MW 6.117 MW
Table 7: Power consumption break-down in several networks.
These numbers are based on information on modern server processors and expected network power draw provided by Eviden. The link power consumption is left constant before considering the power-saving features. The total link count is the number of links on every switch multiplied by the radix, plus the number of NICs. The system is considered idle when the nodes are not performing any computation and thus consume their minimum power (Min). Switch power only comprises the device power draw with no links connected.
On the contrary, under full load (Max), the system consumes approximately 3 MW for all the network topologies. Note that the network’s power consumption is the sum of switch and link power. For instance, the network power for the Megafly is around 260 KW for switches and 499.2 KW for links, for a total of 759.2 KW. In proportion, the network accounts for 18.57 % of power when the system is idling, and up to 13.2 % under full load when all nodes are computing. Compared to RLFTs and 3D-tori in this example, the Megafly achieves a lower network power consumption (the sum of switches and links), both at full load and when idling.
4.3 Power saving parameters, network traffic and performance metrics
In Table 8, we define the power parameters for each power-saving state, such as the power reduction and the time to transition from the power-saving state to the active state ($t_{w}$), and its counterpart ($t_{s}$), derived from the latest EEE specifications. The table also shows the percentage of power that the system devotes to the network upon applying those values to our selected scenario. The calculi for the power contribution of different elements are similar to the one by [2]. Every percentage shown in the table accounts for the contribution to the system energy. The Network column includes the combined power consumption of switches and links. The upper limit for power consumption on links is the difference between the Wake and Fast Wake/Deep Sleep contribution to power under full load.
Port power state Power $t_{w}$ $t_{s}$ Links Network
@@ 4.3 Power saving parameters, network traffic and performance metrics
Wake 24 W - - 12.21 % 8.68 % 18.58 % 13.20 %
Fast Wake 9.6 W -60 % 375 ns 200 ns 4.89 % 3.66 % 20.04 % 13.93 %
Deep Sleep 2.4 W -90 % 4.48 us 2 us 1.22 % 0.92 % 20.87 % 14.32 %
@@
For the evaluation, we have selected VEF traces of two HPC applications (LAMMPS and PATMOS) and two distributed Deep Learning training sessions (MLWF [24] and AlexNet [23]) to study the network behavior in both contexts. For each application, we will first describe the application behavior and the potential for energy saving. First, we will compare how much energy we save with different techniques in comparison to not using any power-saving technique. Non-energy metrics include the degradation that is introduced by power-saving strategies. Mainly, we want to focus on the increase in execution time and latency. Specifically, the latency metric we are exploring is the packet latency, which measures the time from the packet generation to its arrival at the destination.
@@ 4.4 Evaluation for the LAMMPS application
In summary, LAMMPS creates a mix of localized and global communication patterns, with MPI_AllReduce identified as the primary contributor to potential network bottlenecks, especially in large-scale simulations. Being an all-to-all communication operation, MPI_AllReduce will use most paths between nodes. This means that this operation will have the most degradation if all links are in a power-saving state. Figure 6 shows the network efficiency or throughput (normalized against the total network bandwidth) for the complete execution time (2.431 seconds) in the configured system. The total energy consumed is 9.993 MJ by nodes and 1.846 MJ by the network, totaling 11.839 MJ for the whole system. The base execution time (no power-saving mechanisms) is 2.4313 seconds. We can see that after a second of computation, the network activity starts being intermittent with several traffic spikes due to the inter-node collective operations.
@@ 4.4.1 Results with fixed $t_{PDT}$ values
(b) System energy savings.
@@
In general, when using a fixed $t_{PDT}$ value, we can see that the maximum energy savings are close to 10 % when a LAMMPS application is run using our power model and the $t_{PDT}$ values are higher than or equal to 100 µs. With this $t_{PDT}$, the overheads in execution time and latency are also negligible. Fast Wake is more sensitive to smaller $t_{PDT}$ values and achieves better energy savings with smaller $t_{PDT}$ values. Finally, note that with fixed $t_{PDT}$ values larger than 1msms there are barely any energy savings.
@@ 4.4.2 Results using PerfBound and PerfBoundCorrect
The overhead on execution time has a direct impact on the energy consumed, as seen in Figure 8(b). Both PerfBound and PerfBoundCorrect provide similar levels of energy savings, with slightly higher values using PerfBoundCorrect. While PerfBound with Deep Sleep does not lead to energy savings, for thresholds of 1 % and 2 %, PerfBoundCorrect manages to revert that situation. When using a high threshold, however, energy saving does not occur because the calculated $t_{PDT}$ values are too short.
Finally, Figure 8(c) shows that PerfBoundCorrect, in all cases and with either power-saving state, reduces latency significantly. It is clear that, particularly when using Deep Sleep, a circular buffer to control the histogram size produces the highest overhead.
@@ 4.5 Evaluation for the PATMOS application
PATMOS [9] is a Monte Carlo neutron transport code developed at the CEA (France), designed to prototype applications in nuclear safety and radiation shielding.
@@
During each simulation cycle, particles are evenly distributed among threads, with each thread responsible for simulating particle histories and updating the associated scores. By default, static scheduling is used so that the parallelization strategy is deterministic, ensuring reproducibility and simplifying debugging.
@@ 4.5.1 Results with fixed ${t}_{PDT}$ values
The first statistic we will analyze is the execution time, as it is the one that is most noticeable in terms of overhead, and longer execution times have a negative impact on total energy consumption. In our experiments, applying any power-saving technique did not affect this application. The reason behind this is that performance overhead is cumulative with the number of packets that get stalled on output ports. Therefore, as the network is only used at the start and end of the simulation, the effect on execution time is virtually non-existent.
@@
(b) System energy savings.
@@ 4.5.2 Results using PerfBound and PerfBoundCorrect
The execution time with any fixed $t_{PDT}$ value is nearly identical to the configuration without applying any power-saving. PerfBound, focusing on minimizing the overhead on execution time, is no exception. Every execution time with any threshold or histogram management scheme yields the same results.
@@ 4.6 Evaluation for the MLWF application
The MLWF (Machine Learning Weather Forecast) [24] application is developed within the context of the MAELSTROM European project. The trace was collected during a training session using Horovod [32]. The traffic pattern in this application, in particular, and in Deep Learning training sessions in general, usually involves the usage of MPI collective operations for data exchange. Particularly, the most predominant traffic portion comes from AllReduce operations. Usually, every layer of the neural network that is being trained runs a repetitive sequence of Gather-Gatherv-Broadcast-Broadcast until the weights converge, then the AllReduce is invoked to have the information on every node [3].
The network efficiency results for this trace (Figure 12) are shown in logarithmic scale to illustrate that any traffic rate, regardless of the value, provides great difficulty in finding opportunities for power saving.
@@ 4.6.1 Results with fixed ${t}_{PDT}$ values
(c) System energy savings.
@@
The results for energy saving for this application are shown in Figure 13(c). As shown, $t_{PDT}$ values of 10 and 100 µs increase the energy saved in both power-saving states. Lower values provide marginal power-saving values or even increased energy usage in that proportion. Indeed, even if the traffic between the large bursts is irrelevant in terms of bytes transmitted, that communication is not negligible for power saving. However, once we use higher values, the energy reduction begins to diminish. This is due to inactivity, 99 % of periods being within the millisecond range. This is also the reason why using a 1-second $t_{PDT}$ provides more power saving than a millisecond-scale one.
@@ 4.7.1 Results with fixed ${t}_{PDT}$ values
(c) System energy savings.
@@ 5 Outcomes and future work
In this work, we have provided a detailed examination of the implications of power in interconnection networks. First, we explored the simplest technique by using fixed $t_{PDT}$ values, which gave us an upper bound for degradation. Later on, we discussed more sophisticated approaches and solutions that have been proposed, focusing on HPC environments using Ethernet-based networks. Indeed, BXIv3 is only one of several technologies that our proposal is applicable to, such as the upcoming OmniPath versions. After careful examination of our own implementation of the PerfBound technique and the results we gathered, we have found several key aspects that could be improved. As we have mentioned in earlier sections, the results are significant compared to the original PerfBound technique, and prove that giving more breathing room can actually be beneficial to energy efficiency. Furthermore, we also showed that other ostensibly trivial aspects, like histogram management, can have an impact on the energy saved, which was not examined in the original work.
@@
In our humble opinion, our solution is subject to improvements, and several new questions arise. For example, whether providing our dynamic $t_{PDT}$ value calculation with some recency bias would improve the performance on very repetitive patterns.